OpenStudy (anonymous):

Which fraction is between 1 1/2 and 2?

OpenStudy (anonymous):

the options are 1 1/4 1 1/3 2 1/3 1 2/3

OpenStudy (anonymous):

What do you think it is? c:

OpenStudy (texaschic101):

1 1/2 = 1.50 and 2 is just 2 1 1/4 = 1.25 1 1/3 = 1.33 1 2/3 = 1.66 2 1/3 = 2.33 does that help you decide ?

OpenStudy (anonymous):

well kinda but i think it is 1 2/3 or 2 1/3

OpenStudy (texaschic101):

think about 2 1/3....is that number bigger then 2 ?

OpenStudy (anonymous):

2.33 rounds to 2

OpenStudy (texaschic101):

do not round.....

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

it has to be between 1 1/2 and 2....nothing over 2. Your answer is 1 2/3

OpenStudy (anonymous):

ok im about to post another question or should i type it here in this chat @texaschic101 so you could help me